please help!
Translate the following into a quadratic equation, and solve it:
The length of a rectangular garden is three times its width; if the area of the garden is 75 square meters, what are its dimensions?

Answer by rfer(16322)

length=3w
width=w
------------
3w*w=75
3w^2=75
divide both sides by 3
w^2=75/3
w^2=25
square root of 25
w=5
----------------
5*15=75